OpenStudy (emilyjones284):

this is my last one that im having a problem with...ill give a medal (:

OpenStudy (emilyjones284):

\[3\sqrt[3]{81}-2\sqrt[3]{54}\]

OpenStudy (emilyjones284):

simplify.

OpenStudy (anonymous):

\[3 * \sqrt[3]{81} - 2*\sqrt[3]{54} = 3*\sqrt[3]{27*3} - 2*\sqrt[3]{27*2} = 3 *\sqrt[3]{27}*\sqrt[3]{3} - 2*\sqrt[3]{27}*\sqrt[3]{2} \] we know that 3rd root of 27 is equal to 3 (because 3*3*3 = 27) \[3*3*\sqrt[3]{3} - 2*3*\sqrt[3]{2} = 9*\sqrt[3]{3} - 6*\sqrt[3]{2}\]

OpenStudy (emilyjones284):

Thanks (: